4 Simone n4 Find the convergence interval of the entire series A) ]-4,41 B) ]-4,4) C)(-4,4) D) (-4,4[ E) )-1.1/ F))-*..*
Answer #1

D)[-4,4[ is correct option.

ann um E ti 941 tint/ (n+1) anes xn n=1 = Lim noo Ic1+34) = 1X1 </ 4 X-4. n=1 n Ixl <4 at boundary 8 (-1) → converges by Alte